Find three consecutive odd integers such that the sum of the first and the second is 27 less than three times the third
Answer by Greenfinch(383)

Call the numbers N, N+1, N+2
Condition is 2N + 1 + 27 = 3(N + 2)
So 2N +28 = 3N + 6
2N + 22 = 3N
N = 22
Check, Numbers are 22, 23, 24
So 22 + 23 + 27 = 72 = 3 x 24